OpenStudy (anonymous):

 What is the molar mass of Mg(NO3)2? Explain how you calculated this value.

OpenStudy (matt101):

Each atom, molecule, etc has a molecular mass measured in atomic mass units (amu) that is the sum of its protons and neutrons. A mole of an atom, molecule, etc is the number of atoms, molecules, etc you need to get that same number but in grams. Therefore, just add up the masses of each of the atoms in the compound to find its molecular mass/molar mass. 1 Mg + 2 N + 6 O = 1(24.3) + 2(14) + 6(16) = 148.3 amu One molecule of Mg(NO3)2 has a mass of 148.3 amu, so one mole of Mg(NO3)2 has a mass of 148.3 g.